One container holds 10 L of a 25% alcohol solution. How much water should be added to dilute the solution to only 15%?
Answer by addingup(3677)

0.25(10)+0x = 0.15(10+x)
2.5+0x =1.5+0.15x
1 = 0.15x
1/0.15 = x
x = 6.666... L